​the number of ways in which 85 can be written as a sum of two squares

Grade:8

1 Answers

Ritika Das
6 years ago
The number of ways in which 85 can be written as a sum of two squares is by two ways. They are as follows: – 22+92=4+81=85
62+72=36+49=85
Please approve my answer.
Regards.